# SLS-BRD: A system-level approach to seeking generalised feedback Nash equilibria

### v2

@@ 6             Update $s^{p}_{k+1}\in T^{p}(s^{p}_{k},R^{p}(s_{k})\mid L^{p})$;
The procedure in Algorithm 1 belongs to the class of fixed-point methods: Its termination implies that $s^{*}$ is a fixed-point of both $T^{p}$ and $R^{p}$, that is, $s^{*}\in T^{p}(s^{p^{*}},R^{p}(s^{*}))\subseteq T^{p}(s^{p^{*}},s^{-p^{*}})$. In its general form, the conditions (and convergence rates) for these learning dynamics to approach an equilibrium are difficult to establish. In this work, we build upon a fundamental instance from this class of algorithms: The best-response dynamics (BRD). This routine is overviewed in the following.
@@
Let the mapping $BR^{p}:\mathcal{S}^{-p}\rightrightarrows\mathcal{S}^{p}$,
@@
denote the best-response of $p\in\mathcal{P}$ to other players’ strategies. Collectively, $BR(s)\coloneqq BR^{1}(s^{-1})\times\cdots\times BR^{N_{P}}(s^{-N_{P}})\subseteq\mathcal{S}$ is the joint best-response to any given profile $s\in\mathcal{S}$. Then,
Theorem 1.
A strategy profile $s^{*}=(s^{1^{*}},\ldots,s^{N_{P}^{*}})\in\mathcal{S}$ is a GNE for $\mathcal{G}$ if and only if $s^{*}\in BR(s^{*})$ or, equivalently,
$s^{p^{*}}\in BR^{p}(s^{-p^{*}}),\quad\forall p\in\mathcal{P}.$ (5)
The task of computing a Nash equilibrium can thus be translated into searching for a fixed-point of the set-valued mapping $BR:\mathcal{S}\rightrightarrows\mathcal{S}$. The set of GNE solutions for $\mathcal{G}$ is the set of all such fixed-points, $\Omega_{\mathcal{G}}\coloneqq\{s^{*}\in\mathcal{S}:s^{*}\in BR(s^{*})\}$. A natural procedure for GNE seeking thus consists of players adapting their strategies towards best-responses to their rivals’ strategies, which they assume will remain constant. Formally,
@@
given $R^{p}(s_{k})=s_{k}^{-p}$ and a learning rate factor of $\eta\in(0,1)$. This learning dynamics, summarised in Algorithm 2, is known as best-response dynamics or fictitious play.
@@
Output: GNE $s^{*}=(s^{1^{*}},\ldots,s^{N_{P}^{*}})$
@@
2 Initialize $s_{0}=(s_{0}^{1},\ldots,s_{0}^{N_{P}})$ and $k=0$;
@@ 6             Update $s^{p}_{k+1}\in(1{-}\eta)s^{p}_{k}+\eta BR^{p}(s^{-p}_{k})$ ;
After each episode, the strategy profile gets updated as
@@
Theorem 2.
Let $BR:\mathcal{S}\rightrightarrows\mathcal{S}$ be a non-expansive mapping. Then, the best-response dynamics $s_{k+1}=T(s_{k})$ converge monotonically to a GNE solution $s^{*}\in\Omega_{\mathcal{G}}$, that is,
$\lim_{k\to\infty}\inf_{s^{*}\in\Omega_{\mathcal{G}}}\|T(s_{k})-s^{*}\|=0$ (8)
@@
This convergence result stems from fixed-point theory, where the BRD algorithm is interpreted as belonging to the class of averaged (or Mann-Krasnosel‘skii) iteration methods [38, 39]. For non-generalised quadratic problems, we show in Section III-B that the best-response mapping $BR$ is non-expansive under specific conditions on $L^{1},\ldots,L^{N_{P}}$ Importantly, if the best-response mapping is a contraction, then the BRD converge geometrically to a GNE solution $s^{*}\in\Omega_{\mathcal{G}}$, which is unique.
Theorem 3.
Let $BR:\mathcal{S}\rightrightarrows\mathcal{S}$ be $L_{BR}$-Lipschitz, $L_{BR}<1$. Then, from any feasible $s_{0}\in\mathcal{S}$, the best-response dynamics $s_{k+1}=T(s_{k})$ converge to the unique GNE $s^{*}\in\Omega_{\mathcal{G}}$ with rate
$\cfrac{\|s_{k}-s^{*}\|}{\|s_{0}-s^{*}\|}\leq\big{(}(1{-}\eta)+\eta L_{BR}\big{)}^{k}$ (9)
@@
Importantly, these results might not hold in practice whenever the best-response maps, $\{BR^{p}\}_{p\in\mathcal{P}}$, are only approximated (e.g., by solving Eq. (4) numerically). However, such inexact averaged operators are still known to converge under specific assumptions on the accuracy of this approximation [40]. The learning rate $\eta$ also plays a central role in the numerical stability of the BRD algorithm: A careful choice is required to ensure that strategy updates do not escape the feasible set, that is, to ensure that $T(s_{k})\in\mathcal{S}_{\mathcal{G}}\cap\mathcal{S}$ for all $s_{k}\in\mathcal{S}_{\mathcal{G}}\cap\mathcal{S}$. Otherwise, the learning dynamics become undefined and the aforementioned convergence results do not apply. For non-generalised games, $T$ trivially satisfy the constraints for any $\eta\in(0,1)$ and thus a careful design of the learning rate might not be necessary. In such cases, $\eta\to 1$ is the optimal choice if $BR$ is a contraction and the convergence rate in Eq. (9) simplifies to $L_{BR}^{k}$.
@@
II-A Infinite-horizon dynamic games
An infinite-horizon dynamic $N_{P}$-player game, denoted by tuple
@@ II-A Infinite-horizon dynamic games
$\bm{F_{w}}:\bm{u}\mapsto(I-\bm{S_{+}}\bm{A})^{-1}\Big{(}\displaystyle\sum_{p\in\mathcal{P}}\bm{S_{+}}\bm{B}^{p}\bm{u}^{p}+\bm{S_{+}}\bm{w}+\bm{\delta}x_{0}\Big{)},$ (13)
with $\bm{A}:\bm{x}\mapsto(Ax_{t})_{t\in\mathbb{N}}$ and $\bm{B}^{p}:\bm{u}^{p}\mapsto(B^{p}u^{p}_{t})_{t\in\mathbb{N}}$. Because known, the dependency on $x_{0}$ is omitted to simplify notation. Moreover, we assume $\mathrm{E}w_{t}=0$ and $\mathrm{E}(w_{t}w_{t^{\prime}}^{{\mkern-1.5mu\mathsf{T}}})=\delta_{t-t^{\prime}}\Sigma_{w}$, given a covariance matrix $\Sigma_{w}\in\mathbb{S}_{++}^{N_{x}}$, for every $t,t^{\prime}\in\mathbb{N}$. Finally, the sets $\bm{\mathcal{X}}$, $\bm{\mathcal{U}}^{p}$ ($\forall p\in\mathcal{P}$), and $\bm{\mathcal{W}}$ define all permissible state, action, and noise sequences; they take the form
$\bm{\mathcal{X}}$$\coloneqq\{\bm{x}\in\ell_{\infty}^{N_{x}}(\mathbb{N}):x_{t}\in\mathcal{X},\leavevmode\nobreak\ t\in\mathbb{N}\};$
$\bm{\mathcal{U}}^{p}$$\coloneqq\{\bm{u}^{p}\in\ell_{\infty}^{N_{u}^{p}}(\mathbb{N}):u^{p}_{t}\in\mathcal{U}^{p},\leavevmode\nobreak\ t\in\mathbb{N}\};$
$\bm{\mathcal{W}}$$\coloneqq\{\bm{w}\in\ell_{\infty}^{N_{x}}(\mathbb{N}):w_{t}\in\mathcal{W},\leavevmode\nobreak\ t\in\mathbb{N}\},$
@@
In dynamic stochastic games, each player decides on a plan of action $\bm{u}^{p}\in U^{p}(\bm{u}^{-p})$ to minimize its objective functional
@@
defined by cost function $L^{p}:\mathcal{X}\times\mathcal{U}^{1}\times\cdots\times\mathcal{U}^{N_{P}}\to\mathbb{R}$. The mappings $U^{p}:\bm{\mathcal{U}}^{-p}\rightrightarrows\bm{\mathcal{U}}^{p}$ restrict the permissible strategies for each player based on their rivals’ strategies. Under this formulation, the dynamic game $\mathcal{G}_{\infty}$ is stationary and can be interpreted as a static game defined on the appropriate functional spaces. A plan of action $\bm{u}=(\bm{u}^{1},\ldots,\bm{u}^{N_{P}})\in\bm{\mathcal{U}}$ can then be characterised as a generalised Nash Equilibrium solution to $\mathcal{G}_{\infty}$ when no player can improve its objective by unilaterally deviating from this agreed profile. Formally,
@@
A strategy profile $\bm{u}^{*}=(\bm{u}^{1^{*}},\ldots,\bm{u}^{N_{P}^{*}})\in\bm{\mathcal{U}}$ is a generalized Nash equilibrium (GNE) for the game $\mathcal{G}_{\infty}$ if
$J^{p}(\bm{u}^{p^{*}},\bm{u}^{-p^{*}})\leq\min_{\bm{u}^{p}\in U^{p}(\bm{u}^{-p^{*}})}J^{p}(\bm{u}^{p},\bm{u}^{-p^{*}})$ (15)
@@
is not necessarily a singleton and can include non-admissible equilibria. The game is considered unsolvable if $\Omega_{\mathcal{G}_{\infty}}=\emptyset$. The following assumptions are taken for this class of games:
@@
    $U^{p}(\bm{u}^{-p})\coloneqq\{\bm{u}^{p}\in\bm{\mathcal{U}}^{p}:(\bm{u}^{p},\bm{u}^{-p})\in\bm{F_{w}}^{\dagger}(\bm{\mathcal{X}})\cap\bm{\mathcal{U}}_{\mathcal{G}}\},$
    where $\bm{F_{w}}^{\dagger}$ is a left-inverse of the operator in Eq. (13) and
    $\bm{\mathcal{U}}_{\mathcal{G}}=\{\bm{u}\in\textstyle\prod_{p\in\mathcal{P}}\ell_{\infty}^{N_{u}^{p}}(\mathbb{N}):(u^{p}_{t},u^{-p}_{t})\in\mathcal{U}_{\mathcal{G}},\leavevmode\nobreak\ t\in\mathbb{N}\}$
    is a global constraint set shared by all players. The sets $\mathcal{U}^{p}$, $\mathcal{U}_{\mathcal{G}}$, and $\mathcal{X}$ are all non-empty, compact, and convex. Finally, we have that $\bm{F_{w}}^{\dagger}(\bm{\mathcal{X}})\cap\bm{\mathcal{U}}_{\mathcal{G}}\cap(\bm{\mathcal{U}}^{1}\times\cdots\times\bm{\mathcal{U}}^{N_{P}})\neq\emptyset$.
These conditions are analogous to those of Assumption 1: They ensure the existence of GNE solutions to $\mathcal{G}_{\infty}$, that is, $\Omega_{\mathcal{G}_{\infty}}\neq\emptyset$. In this case, the players not only share a global constraint $\bm{\mathcal{U}}_{\mathcal{G}}$, but are also required to ensure that state trajectories lie in a feasible set, $x\in\bm{\mathcal{X}}$, against all possible noise realisations. These constraints often describe operational desiderata and/or limitations concerning the game $\mathcal{G}_{\infty}$. Moreover, note that Assumption 2 implies that the operator $\bm{F_{w}}$ is injective, that is, that distinct control signals cannot produce the same state trajectory. A sufficient condition for this property consists on the matrix $B=[B^{1}\leavevmode\nobreak\ B^{2}\leavevmode\nobreak\ \cdots\leavevmode\nobreak\ B^{N_{P}}]$ being full-rank.
Any equilibrium $\bm{u}^{*}\in\Omega_{\mathcal{G}_{\infty}}$ incorporates an open-loop information pattern: Actions $u_{t}^{*}=(u_{t}^{1^{*}},\ldots,u_{t}^{N_{P}^{*}})$ depend explicitly only on initial $x_{0}\in\mathcal{X}$ and stage $t\in\mathbb{N}$. A plan of action with such representation is undesirable, as players become sensible to noise disturbances and decision errors. Conversely, a state-feedback decision policy, $\bm{u}^{*}=K(\bm{x})$ for some $K:\bm{\mathcal{X}}\to\bm{\mathcal{U}}$, often carry desirable robustness properties. In this sense, a feedback (respectively, open-loop) representation of $\bm{u}^{*}\in\Omega_{\mathcal{G}_{\infty}}$ is said to be strongly (weakly) time consistent [1]. In this work, we investigate state-feedback solutions to the game $\mathcal{G}_{\infty}$.
@@
given a causal operator $\bm{K}^{p}\in\bm{\mathcal{C}}^{p}\subseteq\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{U}}^{p})$ defined by its convolution kernel $\bm{\Phi}^{p}=(\Phi^{p}_{n})_{n\in\mathbb{N}}\in\ell_{1}(\mathbb{N})$. The sets $\{\bm{\mathcal{C}}^{p}\}_{p\in\mathcal{P}}$ describe the operators that satisfy some $\mathcal{G}_{\infty}$-related restrictions (e.g., information patterns incurred by communication, actuation, and sensing delays). In this setting, players do not devise their actions explicitly but rather by designing a state-feedback policy profile $\bm{K}\coloneqq(\bm{K}^{1},\ldots,\bm{K}^{N_{P}})\in\bm{\mathcal{C}}$, with $\bm{\mathcal{C}}=\bm{\mathcal{C}}^{1}\times\cdots\times\bm{\mathcal{C}}^{N_{P}}$. The solution concept that naturally arises is that of a generalized feedback Nash equilibrium.
@@
A policy profile $\bm{K}^{*}=(\bm{K}^{1^{*}},\ldots,\bm{K}^{N_{P}^{*}})$ is a generalized feedback Nash equilibrium (GFNE) for $\mathcal{G}_{\infty}$ if
$J^{p}(\bm{u}^{p^{*}},\bm{u}^{-p^{*}})\leq\min_{\bm{u}^{p}\in U^{p}(\bm{u}^{-p^{*}})}J^{p}(\bm{u}^{p},\bm{u}^{-p^{*}}),$ (17)
where $\bm{u}^{*}\in\mathrm{Ker}(I-\bm{K}^{*}\bm{F_{w}})$, holds for every $p\in\mathcal{P}$.
@@
We consider $\bm{K}^{*}\in\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}$ to be admissible only if it renders the game stable, that is, if the closed-loop evolution
$\bm{x}^{*}=\big{(}I-\bm{S_{+}}(\bm{A}-\textstyle\sum_{p\in\mathcal{P}}\bm{B}^{p}\bm{K}^{p^{*}})\big{)}^{-1}\big{(}\bm{S_{+}}\bm{w}+\bm{\delta}x_{0}\big{)}$
is bounded ($\bm{x}^{*}\in\ell_{\infty}^{N_{x}}$) for all bounded noise ($\bm{w}\in\ell_{\infty}^{N_{x}}$). A policy satisfying this requirement is said to be stabilising. From Assumption 2, we have $\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}\neq\emptyset$ when $\bm{\mathcal{C}}=\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{U}})$. For a more specific $\bm{\mathcal{C}}\subset\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{U}})$, establishing the existence (and, especially, uniqueness) of a solution is demanding [41]. In practice, the set $\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}$ can be constructed from open-loop equilibria $\bm{u}^{*}\in\Omega_{\mathcal{G}_{\infty}}$ by i) parametrising the set of all possible trajectories $\{\bm{x}_{\bm{w}}^{*}=\bm{F_{w}}\bm{u}^{*}\}_{\bm{w}\in\bm{\mathcal{W}}}$, then ii) identifying policies $(\bm{K}^{1^{*}},\ldots,\bm{K}^{N_{P}^{*}})$ that satisfy $\{\bm{u}^{p^{*}}=\bm{K}^{p^{*}}\bm{x}_{\bm{w}}^{*}\}_{\bm{w}\in\bm{\mathcal{W}}}$, $p\in\mathcal{P}$. Highlighting this equivalence, we refer to such $\bm{u}^{*}$ as the open-loop realisation of the closed-loop policy $\bm{K}^{*}$, and vice-versa.

### v3

@@ 6             Update $s^{p}_{k+1}\in T^{p}(s^{p}_{k},R^{p}(s_{k})\mid L^{p})$;
We investigate algorithms for solving $\mathcal{G}$. A direct computation of a GNE is equivalent to solving $N_{P}$ optimisation problems simultaneously, as implied by Definition 1. Such an approach would require players to be coordinated and their objectives to be public. Conversely, we consider adaptive procedures in which the players learn their GNE strategies independently. In this direction, consider that $\mathcal{G}$ admits episodic repetitions, and let $s_{k}\coloneqq(s_{k}^{1},\ldots,s_{k}^{N_{P}})$ be the strategy profile taken by players $\mathcal{P}$ at the $k$-th episode. A prototypical learning routine for equilibrium seeking is outlined in Algorithm 1, with
  • $T^{p}:\mathcal{S}^{p}\times\mathcal{S}^{-p}\rightrightarrows\mathcal{S}^{p}$ describing how the $p$-th player updates its strategy, based on a prediction of its opponents’ next actions, given its individual objective;
  • $R^{p}:\mathcal{S}\rightrightarrows\mathcal{S}^{-p}$ describing how the $p$-th player predicts its opponents’ strategies for the next episode, based on the strategy profile currently being played.
Algorithm 1 belongs to the class of fixed-point methods: Its termination implies that $s^{\star}$ is a fixed-point of both $T^{p}$ and $R^{p}$, that is, $s^{\star}\in T^{p}(s^{p^{\star}},R^{p}(s^{\star}))\subseteq T^{p}(s^{p^{\star}},s^{-p^{\star}})$. In its general form, it is difficult to establish the conditions (and convergence rates) for these learning dynamics to approach an equilibrium. In this work, we build upon a specific yet fundamental instance of this algorithm: The best-response dynamics (BRD). This routine is overviewed in the following.
@@
Let the map $BR^{p}:\mathcal{S}^{-p}\rightrightarrows\mathcal{S}^{p}$,
@@
denote the best-response of $p\in\mathcal{P}$ to other players’ strategies. Collectively, $BR(s)\coloneqq BR^{1}(s^{-1})\times\cdots\times BR^{N_{P}}(s^{-N_{P}})\subseteq\mathcal{S}$ is the joint best-response to any given profile $s\in\mathcal{S}$. From Definition 1, a strategy profile $s^{\star}=(s^{1^{\star}},\ldots,s^{N_{P}^{\star}})\in\mathcal{S}$ is a GNE for $\mathcal{G}$ whenever $s^{\star}\in BR(s^{\star})$ or, equivalently,
$s^{p^{\star}}\in BR^{p}(s^{-p^{\star}}),\quad\forall p\in\mathcal{P}.$ (4)
The task of computing a Nash equilibrium can thus be translated into the search for a fixed-point of the set-valued mapping $BR:\mathcal{S}\rightrightarrows\mathcal{S}$ [7]. The set of GNE solutions for $\mathcal{G}$ is the set of all such fixed-points, $\Omega_{\mathcal{G}}\coloneqq\{s^{\star}\in\mathcal{S}:s^{\star}\in BR(s^{\star})\}$. A natural procedure for GNE seeking consists of players adapting their strategies towards best-responses to their rivals’ strategies, which they assume will remain constant. Formally,
@@
given $R^{p}(s_{k})=s_{k}^{-p}$ and a learning rate factor of $\eta\in(0,1)$. This learning dynamics, summarised in Algorithm 2, is known as (discrete-time) best-response dynamics.
@@
Output: GNE $s^{\star}=(s^{1^{\star}},\ldots,s^{N_{P}^{\star}})$
@@
2 Initialize $s_{0}\coloneqq(s_{0}^{1},\ldots,s_{0}^{N_{P}})$ and $k\coloneqq 0$;
@@ 6             Update $s^{p}_{k+1}\in(1{-}\eta)s^{p}_{k}+\eta BR^{p}(s^{-p}_{k})$ ;
After each episode, the strategy profile is updated to
@@
Lemma 1.
@@
This convergence result stems from fixed-point theory, where the BRD algorithm is interpreted as belonging to the class of averaged (or Krasnosel‘skii-Mann) iteration methods (see [37] for a formal proof). Moreover, if the best-response mapping $BR$ is a contraction then so is $T$ and the BRD must converge geometrically to a GNE $s^{\star}\in\Omega_{\mathcal{G}}$, which is unique [37]:
Lemma 2.
Let $BR:\mathcal{S}\rightrightarrows\mathcal{S}$ be $L_{BR}$-Lipschitz, $L_{BR}<1$. Then, from any feasible $s_{0}\in\mathcal{S}$, the best-response dynamics $s_{k+1}=T(s_{k})$ converge to the unique GNE $s^{\star}\in\Omega_{\mathcal{G}}$ with rate
$\cfrac{\|s_{k}-s^{\star}\|}{\|s_{0}-s^{\star}\|}\leq\big{(}(1{-}\eta)+\eta L_{BR}\big{)}^{k}$ (7)
@@
Importantly, these results might not hold in practice whenever the best-response maps, $\{BR^{p}\}_{p\in\mathcal{P}}$, are only approximated (e.g., by solving Eq. (3) numerically). However, such inexact averaged operators are still known to converge under reasonable assumptions on the accuracy of this approximation [38]. The learning rate $\eta$ plays a central role in the numerical stability of the BRD algorithm: A careful choice is required to ensure that strategy updates do not escape the feasible set, that is, to ensure that $T(s_{k})\in\mathcal{S}_{\mathcal{G}}\cap\mathcal{S}$ for all $s_{k}\in\mathcal{S}_{\mathcal{G}}\cap\mathcal{S}$. The choice of $\eta$ can also ensure convergence in specific games in which $L_{BR}>1$ (see the Supplementary Material). For non-generalised games, $T$ trivially satisfy the constraints for any $\eta\in(0,1)$ and thus a careful design of the learning rate might not be necessary. In such cases, $\eta\to 1$ is the optimal choice if $BR$ is a contraction and the convergence rate in Eq. (7) simplifies to $L_{BR}^{k}$.
@@
2.1 Infinite-horizon dynamic games
A dynamic $N_{P}$-player game, denoted by a tuple
@@ 2.1 Infinite-horizon dynamic games
$\bm{F_{w}}:\bm{u}\mapsto(I-\bm{S_{+}}\bm{A})^{-1}\Big{(}\textstyle\sum_{p\in\mathcal{P}}\bm{S_{+}}\bm{B}^{p}\bm{u}^{p}+\bm{S_{+}}\bm{w}+\bm{\delta}x_{0}\Big{)},$ (11)
with $\bm{A}:\bm{x}\mapsto(Ax_{t})_{t\in\mathbb{N}}$ and $\bm{B}^{p}:\bm{u}^{p}\mapsto(B^{p}u^{p}_{t})_{t\in\mathbb{N}}$. Because known, the dependency on $x_{0}$ is omitted to simplify notation. Moreover, we assume $\mathrm{E}w_{t}=0$ and $\mathrm{E}(w_{t+\tau}w_{t}^{{\mkern-1.5mu\mathsf{T}}})=\delta_{\tau}\Sigma_{w}$, given a covariance matrix $\Sigma_{w}\in\mathbb{S}_{++}^{N_{x}}$, for every $t,\tau\in\mathbb{N}$. Finally, the sets $\bm{\mathcal{X}}$, $\bm{\mathcal{U}}^{p}$ ($\forall p\in\mathcal{P}$), and $\bm{\mathcal{W}}$ define all permissible state, action, and noise sequences; they take the form
$\bm{\mathcal{X}}$$\coloneqq\{\bm{x}\in\ell_{\infty}^{N_{x}}(\mathbb{N}):x_{t}\in\mathcal{X},~{}t\in\mathbb{N}\};$
$\bm{\mathcal{U}}^{p}$$\coloneqq\{\bm{u}^{p}\in\ell_{\infty}^{N_{u}^{p}}(\mathbb{N}):u^{p}_{t}\in\mathcal{U}^{p},~{}t\in\mathbb{N}\};$
$\bm{\mathcal{W}}$$\coloneqq\{\bm{w}\in\ell_{\infty}^{N_{x}}(\mathbb{N}):w_{t}\in\mathcal{W},~{}t\in\mathbb{N}\},$
@@
In infinite-horizon games, each player chooses a plan of action $\bm{u}^{p}\in U^{p}(\bm{u}^{-p})$ to minimize its objective functional
@@
defined by cost function $L^{p}:\mathcal{X}\times\mathcal{U}^{1}\times\cdots\times\mathcal{U}^{N_{P}}\to\mathbb{R}$. The mappings $U^{p}:\bm{\mathcal{U}}^{-p}\rightrightarrows\bm{\mathcal{U}}^{p}$ restrict the permissible actions for each player based on its rivals’ strategies. Under this setup, the dynamic game $\mathcal{G}_{\infty}$ is stationary and can be interpreted as a static game on the appropriate functional spaces. A plan of action $\bm{u}=(\bm{u}^{1},\ldots,\bm{u}^{N_{P}})\in\bm{\mathcal{U}}$ can then be characterised as a GNE solution to $\mathcal{G}_{\infty}$ when no player can improve its objective by unilaterally deviating from this profile. Formally,
@@
A strategy profile $\bm{u}^{\star}=(\bm{u}^{1^{\star}},\ldots,\bm{u}^{N_{P}^{\star}})\in\bm{\mathcal{U}}$ is a generalized Nash equilibrium (GNE) for the game $\mathcal{G}_{\infty}$ if
$J^{p}(\bm{u}^{p^{\star}},\bm{u}^{-p^{\star}})\leq\min_{\bm{u}^{p}\in U^{p}(\bm{u}^{-p^{\star}})}J^{p}(\bm{u}^{p},\bm{u}^{-p^{\star}})$ (13)
@@
is not necessarily a singleton and the game is considered unsolvable if $\Omega_{\mathcal{G}_{\infty}}=\emptyset$. The following assumptions are taken:
@@
    $U^{p}(\bm{u}^{-p})\coloneqq\{\bm{u}^{p}\in\bm{\mathcal{U}}^{p}:(\bm{u}^{p},\bm{u}^{-p})\in\bm{\mathcal{U}}_{\mathcal{G}}\},$
    given the global constraint set
    $\bm{\mathcal{U}}_{\mathcal{G}}=\{\bm{u}\in\ell_{\infty}^{N_{u}}(\mathbb{N}):(F_{w}u)_{t}\in\mathcal{X},~{}u_{t}\in\mathcal{U}_{\mathcal{G}},~{}t\in\mathbb{N}\}.$
    The sets $\mathcal{U}^{p}$, $\mathcal{U}_{\mathcal{G}}$, and $\mathcal{X}$ are all nonempty, compact, and convex. Finally, we have that $\bm{\mathcal{U}}_{\mathcal{G}}\cap(\bm{\mathcal{U}}^{1}\times\cdots\times\bm{\mathcal{U}}^{N_{P}})\neq\emptyset$.
These conditions are analogous to those of Assumption 1: They aim to ensure the existence of GNE solutions to $\mathcal{G}_{\infty}$, that is, $\Omega_{\mathcal{G}_{\infty}}\neq\emptyset$. Here, the shared constraints $\bm{\mathcal{U}}_{\mathcal{G}}$ also require the players to ensure that state trajectories lie in a feasible set ($x\in\bm{\mathcal{X}}$) against all realisations of the noise. These constraints describe operational desiderata and/or limitations in the game.
The equilibria $\bm{u}^{\star}\in\Omega_{\mathcal{G}_{\infty}}$ have an open-loop information pattern: Actions $u_{t}^{\star}=(u_{t}^{1^{\star}},\ldots,u_{t}^{N_{P}^{\star}})$ depend explicitly only on initial state $x_{0}\in\mathcal{X}$ and stage-index $t\in\mathbb{N}$. A plan of action with such representation is undesirable, as players become sensible to noise disturbances and decision errors. Conversely, state-feedback policies $\bm{u}^{\star}=K(\bm{x})$, for some $K:\bm{\mathcal{X}}\to\bm{\mathcal{U}}$, can detect such errors and provide corrective actions. A feedback (respectively, open-loop) representation of $\bm{u}^{\star}\in\Omega_{\mathcal{G}_{\infty}}$ is thus said to be strongly (weakly) time consistent [1]. In this work, we investigate state-feedback solutions to the game $\mathcal{G}_{\infty}$.
@@
given a linear causal operator $\bm{K}^{p}\in\bm{\mathcal{C}}^{p}\subseteq\mathcal{L}(\ell_{\infty}^{N_{x}},\ell_{\infty}^{N_{u}^{p}})$ defined by its convolution kernel $\bm{\Phi}^{p}=(\Phi^{p}_{n})_{n\in\mathbb{N}}\in\ell_{1}(\mathbb{N})$. The sets $\{\bm{\mathcal{C}}^{p}\}_{p\in\mathcal{P}}$ describe the operators that satisfy some $\mathcal{G}_{\infty}$-related restrictions (e.g., information patterns incurred by communication, actuation, and sensing delays). In this setup, players do not plan their actions explicitly but rather by designing a state-feedback policy profile $\bm{K}\coloneqq(\bm{K}^{1},\ldots,\bm{K}^{N_{P}})\in\bm{\mathcal{C}}$, with $\bm{\mathcal{C}}=\bm{\mathcal{C}}^{1}\times\cdots\times\bm{\mathcal{C}}^{N_{P}}$. The solution concept that naturally arises is that of a generalized feedback Nash equilibrium.
@@
A policy profile $\bm{K}^{\star}=(\bm{K}^{1^{\star}},\ldots,\bm{K}^{N_{P}^{\star}})$ is a generalized feedback Nash equilibrium (GFNE) for $\mathcal{G}_{\infty}$ if
$J^{p}(\bm{u}^{p^{\star}},\bm{u}^{-p^{\star}})\leq\min_{\bm{u}^{p}\in U^{p}(\bm{u}^{-p^{\star}})}J^{p}(\bm{u}^{p},\bm{u}^{-p^{\star}}),$ (15)
where $\bm{u}^{\star}\in\mathrm{Ker}(I-\bm{K}^{\star}\bm{F_{w}})$, holds for every $p\in\mathcal{P}$.
@@
We consider $\bm{K}^{\star}\in\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}$ to be admissible only if it renders the game stable, that is, if the closed-loop evolution
$\bm{x}^{\star}=\big{(}I-\bm{S_{+}}(\bm{A}-\textstyle\sum_{p\in\mathcal{P}}\bm{B}^{p}\bm{K}^{p^{\star}})\big{)}^{-1}\big{(}\bm{S_{+}}\bm{w}+\bm{\delta}x_{0}\big{)}$
is bounded ($\bm{x}^{\star}\in\ell_{\infty}^{N_{x}}$) for all bounded noise ($\bm{w}\in\ell_{\infty}^{N_{x}}$). A policy satisfying this requirement is said to be stabilising. From Assumption 2, we have that $\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}\neq\emptyset$ when $\bm{\mathcal{C}}=\bm{\mathcal{U}}^{\bm{\mathcal{X}}}$. In the case of $\bm{\mathcal{C}}\subseteq\mathcal{L}(\ell_{\infty}^{N_{x}},\ell_{\infty}^{N_{u}})$, establishing the existence (and, especially, uniqueness) of a solution is demanding [39]. In practice, the set $\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}$ could be constructed from open-loop equilibria $\bm{u}^{\star}\in\Omega_{\mathcal{G}_{\infty}}$ by i) parametrising the set of all possible trajectories $\{\bm{x}_{\bm{w}}^{\star}=\bm{F_{w}}\bm{u}^{\star}\}_{\bm{w}\in\bm{\mathcal{W}}}$, then ii) identifying policies $(\bm{K}^{1^{\star}},\ldots,\bm{K}^{N_{P}^{\star}})$ that satisfy $\{\bm{u}^{p^{\star}}=\bm{K}^{p^{\star}}\bm{x}_{\bm{w}}^{\star}\}_{\bm{w}\in\bm{\mathcal{W}}}$, $p\in\mathcal{P}$. Highlighting this equivalence, we refer to such $\bm{u}^{\star}$ as an open-loop realisation of the closed-loop policy $\bm{K}^{\star}$, and vice-versa.